OpenStudy (cheska_p):

How do you find the domain of f(p) = sqrt (2 - sqrt(p))?

OpenStudy (swissgirl):

Well we know that we cant have a negative value in the square root So \(2-\sqrt{ p} \geq 0\) So lets solve for p and find our domain or in other words the condition that will make this equation true \[ 2-\sqrt{p} \geq 0\] \[2 \geq \sqrt {p}\] \[4 \geq p\]

OpenStudy (cheska_p):

Thank you so much! This make so much sense

OpenStudy (swissgirl):

np :)